OpenStudy (anonymous):

MEDAL WILL BE AWARDED 2 QUESTIONS ONLY. Some information about the surface area and volume of two similar solids has been given. Find the missing value. Question 1 Solid 1: S.A = 324yd^2 V=4104yd^3 Solid2: S.A = ? V= 152yd^3 Question 2: Solid1 S.A = 48m^2 V= 1280m^3 Solid 2 S.A =147m^2 V=?

OpenStudy (anonymous):

@ganeshie8 @ranga @eliassaab @Preetha @secret66 @Taylor<3sRin @Jonathan1997 @musa19965

OpenStudy (ranga):

Similar solids will have the same Volume to Surface Area ratio: \[\Large \left(\frac{Volume}{Surface~Area}\right)_{solid~1} = ~~~\left(\frac{Volume}{Surface~Area}\right)_{solid~2}\]

OpenStudy (anonymous):

Thank you SO MUCH @ranga